OpenStudy (anonymous):

How to find the vertex (x+4)(5x+2) so confused someone help me

OpenStudy (marigirl):

vertex means the turning point. which means it will occur in the middle of both the x intercepts. would u be able to tell me the two x intercepts?

OpenStudy (marigirl):

in your equation

OpenStudy (anonymous):

I already answered this for you \[(-\frac{ b }{ 2a },f(-\frac{ b }{ 2a }))\] once you foil your binomials you will have Ax^2+Bx+C=0